OpenStudy (anonymous):

When graphing the following equations, which one will be a straight line? A. y = x^2 + 2 B. f(x) = 2/3x + 1 C. y = |x + 4| D. x^ + y^ = 1

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Oh, I thought when dealing with x squared , you will see a curve?

OpenStudy (anonymous):

ya x^2 is a parabola

OpenStudy (anonymous):

it would be B because it is a linear equation

OpenStudy (anonymous):

thats what I think too, but y = |x + 4| is my other choice and so I am confused again

OpenStudy (anonymous):

since it is an absolute value the graph would loook like 

OpenStudy (anonymous):

thus not being a strait line

OpenStudy (anonymous):

do you understand why?

OpenStudy (anonymous):

V" is the result of the negative absolute values becoming positive?

OpenStudy (anonymous):

if B) is straight line is the slope m= 2/3 ? yes or no? so that f(x) = (2/3)x + 1 ? yes ?

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

not a problem